# Determination of the density in the linear elastic wave equation

### v2

@@ 4. S-waves and proof of the main theorem
$u_{\alpha}=A_{S}c_{S}\eta_{\alpha}.$
If we consider the equation (13) as an ODE along a geodesic $\gamma$ with respect to the metric $g_{S}$, we can see that $\eta$ is a parallel vector field along $\gamma$. So one can take $|\eta|^{2}_{g_{S}}=1$. We obtain the following equation for $A_{S}$,
$c_{S}^{-2}\left(2\mu c_{S}\partial_{3}A_{S}+(\partial_{3}\mu)c_{S}A_{S}-\partial_{3}c_{S}\mu A_{S}+\frac{1}{2}\mu g^{\beta\gamma}\frac{\partial g_{\beta\gamma}}{\partial\tau}c_{S}A_{S}\right)=0.$
Similar to the $P$-waves, the solution is
$A_{S}=cJ^{-1/2}c_{S}^{-1/2}\rho^{-1/2}$
with some constant $c$. As in previous section, we take $c=1$. Observe that one can not get any information about $\rho$ in the interior of $\Omega$ from the leading order term of the amplitude at the boundary, which is also pointed out in [20].
Next we determine $a_{13}$. For this, we also calculate
@@
$=$ $\lambda(\partial_{\beta}a_{0\alpha}-\Gamma^{\gamma}_{\alpha\beta}a_{0\gamma})g^{\alpha\beta}-\lambda\Gamma^{\alpha}_{33}a_{0\alpha}g^{33}+2g^{33}\mu(-\Gamma_{33}^{\alpha}a_{0\alpha})$
$=$ $\lambda g^{\alpha\beta}\partial_{\beta}a_{0\alpha}-g^{\alpha\beta}\lambda\Gamma^{\gamma}_{\alpha\beta}a_{0\gamma}+(\lambda+2\mu)c_{S}^{-1}g^{\alpha\beta}\partial_{\beta}c_{S}a_{0\alpha},$
@@
$=$ $-c_{S}^{-1}\mu g^{\alpha\beta}\partial_{\beta}c_{S}a_{0\alpha}+g^{\beta\gamma}\Gamma^{\alpha}_{\beta\gamma}\mu a_{0\alpha}.$
Equating the $\mathcal{O}(\varrho)$-terms in the equation (6) with $i=3$, we obtain
$\begin{split}\rho a_{13}-(\lambda+2\mu)c_{S}^{-2}a_{13}\\ +\mathrm{i}\Big{(}g^{\alpha\beta}\partial_{\beta}(\mu a_{0\alpha})+\lambda g^{\alpha\beta}\partial_{\beta}a_{0\alpha}-\lambda g^{\alpha\beta}\Gamma^{\gamma}_{\alpha\beta}a_{0\gamma}+(\lambda+2\mu)c_{S}^{-1}g^{\alpha\beta}\partial_{\beta}c_{S}a_{0\alpha}\\ +c_{S}^{-1}\mu g^{\alpha\beta}\partial_{\beta}c_{S}a_{0\alpha}-g^{\beta\gamma}\Gamma^{\alpha}_{\beta\gamma}\mu a_{0\alpha}\Big{)}=0,\end{split}$
@@
$\begin{split}\rho a_{13}-(\lambda+2\mu)c_{S}^{-2}a_{13}\\ +\mathrm{i}\Big{(}(\lambda+\mu)g^{\alpha\beta}\partial_{\beta}a_{0\alpha}+g^{\alpha\beta}(\partial_{\beta}\mu)a_{0\alpha}+(\lambda+3\mu)c_{S}^{-1}g^{\alpha\beta}\partial_{\beta}c_{S}a_{0\alpha}-(\lambda+\mu)g^{\beta\gamma}\Gamma^{\alpha}_{\beta\gamma}a_{0\alpha}\Big{)}=0.\end{split}$
@@
$\begin{split}a_{13}=&\frac{\mathrm{i}}{c_{P}^{2}c_{S}^{-2}-1}\left(\frac{\lambda+\mu}{\rho}g^{\alpha\beta}\partial_{\beta}a_{0\alpha}+g^{\alpha\beta}\frac{\partial_{\beta}\mu}{\rho}a_{0\alpha}+\frac{\lambda+3\mu}{\rho}c_{S}^{-1}g^{\alpha\beta}\partial_{\beta}c_{S}a_{0\alpha}-\frac{\lambda+\mu}{\rho}g^{\beta\gamma}\Gamma^{\alpha}_{\beta\gamma}a_{0\alpha}\right)\\ =&\frac{\mathrm{i}}{c_{P}^{2}c_{S}^{-2}-1}\Bigg{(}(c_{P}^{2}-c_{S}^{2})g^{\beta\gamma}\partial_{\beta}a_{0\gamma}+g^{\alpha\beta}\partial_{\beta}c_{S}^{2}a_{0\gamma}+2g^{\beta\gamma}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}a_{0\gamma}+(c_{P}^{2}+c_{S}^{2})c_{S}^{-1}g^{\beta\gamma}\partial_{\beta}c_{S}a_{0\gamma}\\ &\quad\quad\quad\quad-(c_{P}^{2}-c_{S}^{2})g^{\beta\gamma}\Gamma^{\alpha}_{\beta\gamma}a_{0\alpha}\Bigg{)}.\end{split}$
@@
$\partial_{\alpha}a_{13}=$ $\frac{\mathrm{i}}{c_{P}^{2}c_{S}^{-2}-1}\Bigg{(}(c_{P}^{2}-c_{S}^{2})g^{\beta\gamma}\partial_{\alpha}\partial_{\beta}a_{0\gamma}+\partial_{\alpha}(c_{P}^{2}-c_{S}^{2})g^{\beta\gamma}\partial_{\beta}a_{0\gamma}+(c_{P}^{2}-c_{S}^{2})\partial_{\alpha}g^{\beta\gamma}\partial_{\beta}a_{0\gamma}$
$\quad\quad+g^{\beta\gamma}(\partial_{\beta}c_{S}^{2})\partial_{\alpha}a_{0\gamma}+2g^{\beta\gamma}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}\partial_{\alpha}a_{0\gamma}+2\partial_{\alpha}g^{\beta\gamma}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}a_{0\gamma}$
$\quad\quad+2g^{\beta\gamma}\partial_{\alpha}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}a_{0\gamma}+2g^{\beta\gamma}c_{S}^{2}\partial_{\alpha}\partial_{\beta}\log\sqrt{\rho}a_{0\gamma}$
$\quad\quad+(c_{P}^{2}+c_{S}^{2})c_{S}^{-1}g^{\beta\gamma}\partial_{\beta}c_{S}\partial_{\alpha}a_{0\gamma}-(c_{P}^{2}+c_{S}^{2})g^{\beta\gamma}\Gamma^{\delta}_{\beta\gamma}\partial_{\alpha}a_{0\delta}$
$\quad\quad-\frac{1}{c_{S}^{2}}\frac{c_{S}^{2}\partial_{\alpha}c_{P}^{2}+c_{P}^{2}\partial_{\alpha}c_{S}^{2}}{c_{P}^{2}-c_{S}^{2}}(c_{P}^{2}-c_{S}^{2})g^{\beta\gamma}\partial_{\beta}a_{0\gamma}-\frac{2}{c_{S}^{2}}\frac{c_{S}^{2}\partial_{\alpha}c_{P}^{2}+c_{P}^{2}\partial_{\alpha}c_{S}^{2}}{c_{P}^{2}-c_{S}^{2}}g^{\beta\gamma}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}a_{0\gamma}\Bigg{)}$
@@
$=$ $\frac{\mathrm{i}}{c_{P}^{2}c_{S}^{-2}-1}\Bigg{(}(c_{P}^{2}-c_{S}^{2})g^{\beta\gamma}\partial_{\alpha}\partial_{\beta}a_{0\gamma}-\frac{c_{P}^{2}+c_{S}^{2}}{c_{S}^{2}}\partial_{\alpha}(c_{S}^{2})g^{\beta\gamma}\partial_{\beta}a_{0\gamma}+(c_{P}^{2}-c_{S}^{2})\partial_{\alpha}g^{\beta\gamma}\partial_{\beta}a_{0\gamma}$
$\quad\quad+2g^{\beta\gamma}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}\partial_{\alpha}a_{0\gamma}+2\partial_{\alpha}g^{\beta\gamma}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}a_{0\gamma}$
$\quad\quad+2g^{\beta\gamma}c_{S}^{2}\partial_{\alpha}\partial_{\beta}\log\sqrt{\rho}a_{0\gamma}$
$\quad\quad+(c_{P}^{2}+3c_{S}^{2})c_{S}^{-1}g^{\beta\gamma}\partial_{\beta}c_{S}\partial_{\alpha}a_{0\gamma}-(c_{P}^{2}-c_{S}^{2})g^{\beta\gamma}\Gamma^{\delta}_{\beta\gamma}\partial_{\alpha}a_{0\delta}$
$\quad\quad+2g^{\beta\gamma}\partial_{\alpha}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}a_{0\gamma}-2\frac{c_{S}^{2}\partial_{\alpha}c_{P}^{2}+c_{P}^{2}\partial_{\alpha}c_{S}^{2}}{c_{P}^{2}-c_{S}^{2}}g^{\beta\gamma}\partial_{\beta}\log\sqrt{\rho}a_{0\gamma}\Bigg{)}$
@@
$=$ $\partial_{\alpha}(\lambda a_{13}g^{33})+\lambda a_{13}g^{33}g^{\beta\gamma}\partial_{\beta}g_{\alpha\gamma}+\partial_{3}(\mu a_{1\alpha})g^{33}$
$=$ $\partial_{\alpha}(\lambda a_{13}c_{S}^{-2})+\lambda a_{13}c_{S}^{-2}g^{\beta\gamma}\partial_{\beta}g_{\alpha\gamma}+\partial_{3}(\mu a_{1\alpha})c_{S}^{-2},$
@@
$=$ $\mu c_{S}^{-2}\partial_{3}a_{1\alpha}+\mu c_{S}^{-2}\partial_{\alpha}a_{13}-2\mu c_{S}^{-2}(\frac{1}{2}g^{\beta\gamma}\frac{\partial g_{\alpha\gamma}}{\partial\tau}a_{1\beta}+c_{S}^{-1}\partial_{\alpha}c_{S}a_{13})$
$=$ $\mu c_{S}^{-2}\partial_{3}a_{1\alpha}-\mu c_{S}^{-2}g^{\beta\gamma}\frac{\partial g_{\alpha\gamma}}{\partial\tau}a_{1\beta}+\mu c_{S}^{-2}\partial_{\alpha}a_{13}-2\mu c_{S}^{-3}\partial_{\alpha}c_{S}a_{13},$
@@
$=$ $g^{jm}\Gamma^{\beta}_{jm}\lambda a_{13}g^{33}g_{\alpha\beta}+(g^{33}\Gamma^{3}_{33}+g^{\beta\gamma}\Gamma^{3}_{\beta\gamma})\mu a_{1\alpha}$
$=$ $-\lambda c_{S}^{-3}\partial_{\alpha}c_{S}a_{13}+c_{S}^{-2}g^{\gamma\delta}g_{\alpha\beta}\Gamma^{\beta}_{\gamma\delta}\lambda a_{13}+c_{S}^{-3}\partial_{3}c_{S}\mu a_{1\alpha}-\frac{1}{2}g^{\beta\gamma}\frac{\partial g_{\beta\gamma}}{\partial\tau}c_{S}^{-2}\mu a_{1\alpha},$
@@
$=$ $g^{\beta\gamma}\partial_{\beta}(\lambda g^{33}(-\Gamma_{33}^{\delta}a_{0\delta})g_{\alpha\gamma}+\lambda(\partial_{\varepsilon}a_{0\delta}-\Gamma_{\varepsilon\delta}^{\eta}a_{0\eta})g^{\delta\varepsilon}g_{\alpha\gamma})$
@@
$=$ $g^{\beta\gamma}\partial_{\beta}(\lambda(\partial_{\varepsilon}a_{0\delta})g^{\delta\varepsilon}g_{\alpha\gamma})-g^{\beta\gamma}\partial_{\beta}(\lambda g^{33}\Gamma^{\delta}_{33}a_{0\delta}g_{\alpha\gamma}+\lambda\Gamma^{\eta}_{\varepsilon\delta}a_{0\eta}g^{\delta\varepsilon}g_{\alpha\gamma})$
$+g^{33}\partial_{3}(\mu\partial_{3}a_{0\alpha})+g^{\beta\gamma}\partial_{\beta}(\mu\partial_{\gamma}a_{0\alpha})+g^{\beta\gamma}\partial_{\gamma}(\mu\partial_{\alpha}a_{0\beta})-2g^{33}\partial_{3}(\mu\Gamma^{\beta}_{\alpha 3}a_{0\beta})-2g^{\delta\gamma}\partial_{\delta}(\mu\Gamma^{\beta}_{\alpha\gamma}a_{0\beta})$
$=$ $\partial_{\alpha}\lambda g^{\delta\varepsilon}\partial_{\varepsilon}a_{0\delta}+\lambda g^{\delta\varepsilon}\partial_{\alpha}\partial_{\varepsilon}a_{0\delta}+\lambda\partial_{\varepsilon}a_{0\delta}\partial_{\alpha}g^{\delta\varepsilon}+\lambda g^{\beta\gamma}g^{\delta\varepsilon}\partial_{\beta}g_{\alpha\gamma}\partial_{\varepsilon}a_{0\delta}$
@@
$+c_{S}^{-2}\partial_{3}\mu\partial_{3}a_{0\alpha}+c_{S}^{-2}\mu\partial^{2}_{33}a_{0\alpha}+g^{\beta\gamma}\partial_{\beta}\mu\partial_{\gamma}a_{0\alpha}+\mu g^{\beta\gamma}\partial_{\beta}\partial_{\gamma}a_{0\alpha}+g^{\beta\gamma}\partial_{\gamma}\mu\partial_{\alpha}a_{0\beta}+\mu g^{\beta\gamma}\partial_{\alpha}\partial_{\gamma}a_{0\beta}$
$-2c_{S}^{-2}\mu\Gamma^{\beta}_{\alpha 3}\partial_{3}a_{0\beta}-2g^{\delta\gamma}\mu\Gamma^{\beta}_{\alpha\gamma}\partial_{\delta}a_{0\beta}-2c_{S}^{-2}\partial_{3}\mu\Gamma^{\beta}_{\alpha 3}a_{0\beta}-2g^{\delta\gamma}\partial_{\delta}\mu\Gamma^{\beta}_{\alpha\gamma}a_{0\beta},$
@@
$=$ $(\Gamma^{3}_{3\alpha}+\Gamma^{\delta}_{\alpha\delta})\lambda g^{\beta\gamma}\partial_{\gamma}a_{0\beta}+g^{33}\Gamma^{\beta}_{\alpha 3}\mu\partial_{3}a_{0\beta}+g^{\gamma\delta}\Gamma^{\beta}_{\alpha\gamma}\mu\partial_{\delta}a_{0\beta}+g^{\beta\gamma}\mu\Gamma^{3}_{\alpha\gamma}\partial_{3}a_{0\beta}+g^{\beta\gamma}\Gamma^{\delta}_{\alpha\gamma}\mu\partial_{\delta}a_{0\beta}$
$+\sqrt{\rho}\times\text{terms that do not depend on $\rho$}$
$=$ $(\Gamma^{3}_{3\alpha}+\Gamma^{\delta}_{\alpha\delta})\lambda g^{\beta\gamma}\partial_{\gamma}a_{0\beta}+g^{\gamma\delta}\Gamma^{\beta}_{\alpha\gamma}\mu\partial_{\delta}a_{0\beta}+g^{\beta\gamma}\Gamma^{\delta}_{\alpha\gamma}\mu\partial_{\delta}a_{0\beta}+\sqrt{\rho}\times\text{terms that do notdepend on $\rho$},$

### v3

@@ Appendix A Proof of lemmas
$a_{\alpha}=A_{S}c_{S}\eta_{\alpha}.$
If we consider the equation (18) as an ODE along a geodesic $\gamma$ with respect to the metric $g_{S}$, we can see that $\eta$ is a parallel cotangent vector field along $\gamma$. So one can take $|\eta|^{2}_{g_{S}}=c_{S}^{2}g^{ij}\eta_{i}\eta_{j}=c_{S}^{2}g^{\alpha\beta}\eta_{\alpha}\eta_{\beta}=1$. We obtain the following equation for $A_{S}$,
$c_{S}^{-2}\left(2\mu c_{S}\partial_{3}A_{S}+(\partial_{3}\mu)c_{S}A_{S}-\partial_{3}c_{S}\mu A_{S}+\frac{1}{2}\mu g^{\beta\varepsilon}\frac{\partial g_{\beta\varepsilon}}{\partial\tau}c_{S}A_{S}\right)=0.$
Similar to the proof of Lemma 1, by solving the above ODE, we prove the lemma. ∎
Proof of Lemma 4.
First we determine $a_{13}$. For this, we calculate
@@
$=$ $\lambda(\partial_{\beta}a_{0\alpha}-\Gamma^{\varepsilon}_{\alpha\beta}a_{0\varepsilon})g^{\alpha\beta}-\lambda\Gamma^{\alpha}_{33}a_{0\alpha}g^{33}+2g^{33}\mu(-\Gamma_{33}^{\alpha}a_{0\alpha})$
$=$ $\lambda g^{\alpha\beta}\partial_{\beta}a_{0\alpha}-g^{\alpha\beta}\lambda\Gamma^{\varepsilon}_{\alpha\beta}a_{0\varepsilon}+(\lambda+2\mu)c_{S}^{-1}g^{\alpha\beta}\partial_{\beta}c_{S}a_{0\alpha},$
@@
$=$ $-c_{S}^{-1}\mu g^{\alpha\beta}\partial_{\beta}c_{S}a_{0\alpha}+g^{\beta\varepsilon}\Gamma^{\alpha}_{\beta\varepsilon}\mu a_{0\alpha}.$
Using equation (9) with $i=3$, we obtain
$\begin{split}\rho a_{13}-(\lambda+2\mu)c_{S}^{-2}a_{13}\\ +\mathrm{i}\Big(g^{\alpha\beta}\partial_{\beta}(\mu a_{0\alpha})+\lambda g^{\alpha\beta}\partial_{\beta}a_{0\alpha}-\lambda g^{\alpha\beta}\Gamma^{\varepsilon}_{\alpha\beta}a_{0\varepsilon}+(\lambda+2\mu)c_{S}^{-1}g^{\alpha\beta}\partial_{\beta}c_{S}a_{0\alpha}\\ +c_{S}^{-1}\mu g^{\alpha\beta}\partial_{\beta}c_{S}a_{0\alpha}-g^{\beta\varepsilon}\Gamma^{\alpha}_{\beta\varepsilon}\mu a_{0\alpha}\Big)=0,\end{split}$
@@
$\begin{split}\rho a_{13}-(\lambda+2\mu)c_{S}^{-2}a_{13}\\ +\mathrm{i}\Big((\lambda+\mu)g^{\alpha\beta}\partial_{\beta}a_{0\alpha}+g^{\alpha\beta}(\partial_{\beta}\mu)a_{0\alpha}+(\lambda+3\mu)c_{S}^{-1}g^{\alpha\beta}\partial_{\beta}c_{S}a_{0\alpha}-(\lambda+\mu)g^{\beta\varepsilon}\Gamma^{\alpha}_{\beta\varepsilon}a_{0\alpha}\Big)=0.\end{split}$
@@
$\begin{split}a_{13}=&\frac{\mathrm{i}}{c_{P}^{2}c_{S}^{-2}-1}\left(\frac{\lambda+\mu}{\rho}g^{\alpha\beta}\partial_{\beta}a_{0\alpha}+g^{\alpha\beta}\frac{\partial_{\beta}\mu}{\rho}a_{0\alpha}+\frac{\lambda+3\mu}{\rho}c_{S}^{-1}g^{\alpha\beta}\partial_{\beta}c_{S}a_{0\alpha}-\frac{\lambda+\mu}{\rho}g^{\beta\varepsilon}\Gamma^{\alpha}_{\beta\varepsilon}a_{0\alpha}\right)\\ =&\frac{\mathrm{i}}{c_{P}^{2}c_{S}^{-2}-1}\Bigg((c_{P}^{2}-c_{S}^{2})g^{\beta\varepsilon}\partial_{\beta}a_{0\varepsilon}+g^{\alpha\beta}\partial_{\beta}c_{S}^{2}a_{0\varepsilon}+2g^{\beta\varepsilon}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}a_{0\varepsilon}+(c_{P}^{2}+c_{S}^{2})c_{S}^{-1}g^{\beta\varepsilon}\partial_{\beta}c_{S}a_{0\varepsilon}\\ &\quad\quad\quad\quad-(c_{P}^{2}-c_{S}^{2})g^{\beta\varepsilon}\Gamma^{\alpha}_{\beta\varepsilon}a_{0\alpha}\Bigg).\end{split}$
@@
$\partial_{\alpha}a_{13}=$ $\frac{\mathrm{i}}{c_{P}^{2}c_{S}^{-2}-1}\Bigg((c_{P}^{2}-c_{S}^{2})g^{\beta\varepsilon}\partial_{\alpha}\partial_{\beta}a_{0\varepsilon}+\partial_{\alpha}(c_{P}^{2}-c_{S}^{2})g^{\beta\varepsilon}\partial_{\beta}a_{0\varepsilon}+(c_{P}^{2}-c_{S}^{2})\partial_{\alpha}g^{\beta\varepsilon}\partial_{\beta}a_{0\varepsilon}$
$\quad\quad+g^{\beta\varepsilon}(\partial_{\beta}c_{S}^{2})\partial_{\alpha}a_{0\varepsilon}+2g^{\beta\varepsilon}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}\partial_{\alpha}a_{0\varepsilon}+2\partial_{\alpha}g^{\beta\varepsilon}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}a_{0\varepsilon}$
$\quad\quad+2g^{\beta\varepsilon}\partial_{\alpha}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}a_{0\varepsilon}+2g^{\beta\varepsilon}c_{S}^{2}\partial_{\alpha}\partial_{\beta}\log\sqrt{\rho}a_{0\varepsilon}$
$\quad\quad+(c_{P}^{2}+c_{S}^{2})c_{S}^{-1}g^{\beta\varepsilon}\partial_{\beta}c_{S}\partial_{\alpha}a_{0\varepsilon}-(c_{P}^{2}+c_{S}^{2})g^{\beta\varepsilon}\Gamma^{\delta}_{\beta\varepsilon}\partial_{\alpha}a_{0\delta}$
$\quad\quad-\frac{1}{c_{S}^{2}}\frac{c_{S}^{2}\partial_{\alpha}c_{P}^{2}+c_{P}^{2}\partial_{\alpha}c_{S}^{2}}{c_{P}^{2}-c_{S}^{2}}(c_{P}^{2}-c_{S}^{2})g^{\beta\varepsilon}\partial_{\beta}a_{0\varepsilon}-\frac{2}{c_{S}^{2}}\frac{c_{S}^{2}\partial_{\alpha}c_{P}^{2}+c_{P}^{2}\partial_{\alpha}c_{S}^{2}}{c_{P}^{2}-c_{S}^{2}}g^{\beta\varepsilon}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}a_{0\varepsilon}\Bigg)$
@@
$=$ $\frac{\mathrm{i}}{c_{P}^{2}c_{S}^{-2}-1}\Bigg((c_{P}^{2}-c_{S}^{2})g^{\beta\varepsilon}\partial_{\alpha}\partial_{\beta}a_{0\varepsilon}-\frac{c_{P}^{2}+c_{S}^{2}}{c_{S}^{2}}\partial_{\alpha}(c_{S}^{2})g^{\beta\varepsilon}\partial_{\beta}a_{0\varepsilon}+(c_{P}^{2}-c_{S}^{2})\partial_{\alpha}g^{\beta\varepsilon}\partial_{\beta}a_{0\varepsilon}$
$\quad\quad+2g^{\beta\varepsilon}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}\partial_{\alpha}a_{0\varepsilon}+2\partial_{\alpha}g^{\beta\varepsilon}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}a_{0\varepsilon}$
$\quad\quad+2g^{\beta\varepsilon}c_{S}^{2}\partial_{\alpha}\partial_{\beta}\log\sqrt{\rho}a_{0\varepsilon}$
$\quad\quad+(c_{P}^{2}+3c_{S}^{2})c_{S}^{-1}g^{\beta\varepsilon}\partial_{\beta}c_{S}\partial_{\alpha}a_{0\varepsilon}-(c_{P}^{2}-c_{S}^{2})g^{\beta\varepsilon}\Gamma^{\delta}_{\beta\varepsilon}\partial_{\alpha}a_{0\delta}$
$\quad\quad+2g^{\beta\varepsilon}\partial_{\alpha}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}a_{0\varepsilon}-2\frac{c_{S}^{2}\partial_{\alpha}c_{P}^{2}+c_{P}^{2}\partial_{\alpha}c_{S}^{2}}{c_{P}^{2}-c_{S}^{2}}g^{\beta\varepsilon}\partial_{\beta}\log\sqrt{\rho}a_{0\varepsilon}\Bigg)$
@@
$=$ $\partial_{\alpha}(\lambda a_{13}g^{33})+\lambda a_{13}g^{33}g^{\beta\varepsilon}\partial_{\beta}g_{\alpha\varepsilon}+\partial_{3}(\mu a_{1\alpha})g^{33}$
$=$ $\partial_{\alpha}(\lambda a_{13}c_{S}^{-2})+\lambda a_{13}c_{S}^{-2}g^{\beta\varepsilon}\partial_{\beta}g_{\alpha\varepsilon}+\partial_{3}(\mu a_{1\alpha})c_{S}^{-2},$
@@
$=$ $\mu c_{S}^{-2}\partial_{3}a_{1\alpha}+\mu c_{S}^{-2}\partial_{\alpha}a_{13}-2\mu c_{S}^{-2}(\frac{1}{2}g^{\beta\varepsilon}\frac{\partial g_{\alpha\varepsilon}}{\partial\tau}a_{1\beta}+c_{S}^{-1}\partial_{\alpha}c_{S}a_{13})$
$=$ $\mu c_{S}^{-2}\partial_{3}a_{1\alpha}-\mu c_{S}^{-2}g^{\beta\varepsilon}\frac{\partial g_{\alpha\varepsilon}}{\partial\tau}a_{1\beta}+\mu c_{S}^{-2}\partial_{\alpha}a_{13}-2\mu c_{S}^{-3}\partial_{\alpha}c_{S}a_{13},$
@@
$=$ $g^{jm}\Gamma^{\beta}_{jm}\lambda a_{13}g^{33}g_{\alpha\beta}+(g^{33}\Gamma^{3}_{33}+g^{\beta\varepsilon}\Gamma^{3}_{\beta\varepsilon})\mu a_{1\alpha}$
$=$ $-\lambda c_{S}^{-3}\partial_{\alpha}c_{S}a_{13}+c_{S}^{-2}g^{\varepsilon\delta}g_{\alpha\beta}\varepsilon^{\beta}_{\varepsilon\delta}\lambda a_{13}+c_{S}^{-3}\partial_{3}c_{S}\mu a_{1\alpha}-\frac{1}{2}g^{\beta\varepsilon}\frac{\partial g_{\beta\varepsilon}}{\partial\tau}c_{S}^{-2}\mu a_{1\alpha},$
@@
$=$ $g^{\beta\varepsilon}\partial_{\beta}(\lambda g^{33}(-\Gamma_{33}^{\delta}a_{0\delta})g_{\alpha\varepsilon}+\lambda(\partial_{\varepsilon}a_{0\delta}-\Gamma_{\kappa\delta}^{\eta}a_{0\eta})g^{\delta\kappa}g_{\alpha\varepsilon})$
@@
$=$ $g^{\beta\varepsilon}\partial_{\beta}(\lambda(\partial_{\kappa}a_{0\delta})g^{\delta\kappa}g_{\alpha\varepsilon})-g^{\beta\varepsilon}\partial_{\beta}(\lambda g^{33}\Gamma^{\delta}_{33}a_{0\delta}g_{\alpha\gamma}+\lambda\Gamma^{\eta}_{\kappa\delta}a_{0\eta}g^{\delta\kappa}g_{\alpha\varepsilon})$
$+g^{33}\partial_{3}(\mu\partial_{3}a_{0\alpha})+g^{\beta\varepsilon}\partial_{\beta}(\mu\partial_{\varepsilon}a_{0\alpha})+g^{\beta\varepsilon}\partial_{\varepsilon}(\mu\partial_{\alpha}a_{0\beta})-2g^{33}\partial_{3}(\mu\Gamma^{\beta}_{\alpha 3}a_{0\beta})-2g^{\delta\varepsilon}\partial_{\delta}(\mu\Gamma^{\beta}_{\alpha\varepsilon}a_{0\beta})$
$=$ $\partial_{\alpha}\lambda g^{\delta\varepsilon}\partial_{\varepsilon}a_{0\delta}+\lambda g^{\delta\varepsilon}\partial_{\alpha}\partial_{\varepsilon}a_{0\delta}+\lambda\partial_{\varepsilon}a_{0\delta}\partial_{\alpha}g^{\delta\varepsilon}+\lambda g^{\beta\kappa}g^{\delta\varepsilon}\partial_{\beta}g_{\alpha\kappa}\partial_{\varepsilon}a_{0\delta}$
@@
$+c_{S}^{-2}\partial_{3}\mu\partial_{3}a_{0\alpha}+c_{S}^{-2}\mu\partial^{2}_{33}a_{0\alpha}+g^{\beta\varepsilon}\partial_{\beta}\mu\partial_{\varepsilon}a_{0\alpha}+\mu g^{\beta\varepsilon}\partial_{\beta}\partial_{\varepsilon}a_{0\alpha}+g^{\beta\varepsilon}\partial_{\varepsilon}\mu\partial_{\alpha}a_{0\beta}+\mu g^{\beta\varepsilon}\partial_{\alpha}\partial_{\varepsilon}a_{0\beta}$
$-2c_{S}^{-2}\mu\Gamma^{\beta}_{\alpha 3}\partial_{3}a_{0\beta}-2g^{\delta\varepsilon}\mu\Gamma^{\beta}_{\alpha\varepsilon}\partial_{\delta}a_{0\beta}-2c_{S}^{-2}\partial_{3}\mu\Gamma^{\beta}_{\alpha 3}a_{0\beta}-2g^{\delta\varepsilon}\partial_{\delta}\mu\Gamma^{\beta}_{\alpha\varepsilon}a_{0\beta},$
@@
$=$ $(\Gamma^{3}_{3\alpha}+\Gamma^{\delta}_{\alpha\delta})\lambda g^{\beta\varepsilon}\partial_{\varepsilon}a_{0\beta}+g^{33}\Gamma^{\beta}_{\alpha 3}\mu\partial_{3}a_{0\beta}+g^{\varepsilon\delta}\Gamma^{\beta}_{\alpha\varepsilon}\mu\partial_{\delta}a_{0\beta}+g^{\beta\varepsilon}\mu\Gamma^{3}_{\alpha\varepsilon}\partial_{3}a_{0\beta}+g^{\beta\varepsilon}\Gamma^{\delta}_{\alpha\varepsilon}\mu\partial_{\delta}a_{0\beta}$
[…]
$=$ $(\Gamma^{3}_{3\alpha}+\Gamma^{\delta}_{\alpha\delta})\lambda g^{\beta\varepsilon}\partial_{\varepsilon}a_{0\beta}+g^{\varepsilon\delta}\Gamma^{\beta}_{\alpha\varepsilon}\mu\partial_{\delta}a_{0\beta}+g^{\beta\varepsilon}\Gamma^{\delta}_{\alpha\varepsilon}\mu\partial_{\delta}a_{0\beta}+\sqrt{\rho}\times\text{terms that do not depend on $\rho$},$